I have been trying to find the answer to this equation for hours. It is a substitution method equation. Here it is, y=2x-3 x-y=-4. Please help! I have gotten multiple answers like x=1 and y=5 but my teacher says that isn't right and i also got x=4 and y=8 but that also isn't right. Please can you help me?
Answer by ankor@dixie-net.com(22740)

y = 2x - 3
x - y = -4
Use the 1st equation for substitution, replace y with (2x-3) in the 2nd equation
x - (2x-3) = -4
Removing the brackets changes the sign of 3
x - 2x + 3 = -4
-x = -4 - 3
-x = -7
Multiply both sides by -1
x = 7
:
Find y using the 1st equation, replace x with 7
y = 2(7) - 3
y = 14 - 3
y = 11
:
:
Check these solution in the 2nd equation
x - y = -4
7 - 11 = -4